In order to conserve fuel you modify your car by removing 100 kg of excess weight. How much energy is saved each time you accelerate from 0 to 100 km/hr?

Please explain and show your work.
Answer #1

According to the given problem,

Let the mass of car be M,

Here as the car is accelrating K.E is the only energy applicable.

And as the car is starting from rest intial K.E is always zero

The energy saved is,

E = K.E1 - K.E2

= 1/2Mv2 - 1/2(M-100)v2

= 1/2*100*v2 where v is = 100km/hr = 27.78 m/s

= 0.5*100*27.782.

= 38.58 = 38.6 kJ. [saved energy]

I hope you understood the problem, If yes rate me!! or else comment for a better solution.